How do I solve this? Sally can paint a house in 4 hours. It takes John 6 hours to paint the same house. How long does it take them together?
I am not sure how to set this up.
Thanks!
Lisa Comer

Found 3 solutions by jim_thompson5910, drk, stanbon:
Answer by jim_thompson5910(35256)

See this previous solution.

Answer by drk(1908)

This is job /time. We want the job over time for Sally and for John. We want the time together, x. So we have
(1/4)x + (1/6)x = 1
step 1 - multiply by 12 to get
3x + 2x = 12
5x = 12
x = 12/5.
2.4 hours.

Answer by stanbon(75887)

How do I solve this? Sally can paint a house in 4 hours. It takes John 6 hours to paint the same house. How long does it take them together?
------------------------------------
Sally's rate: 1/4 job/hr
John's rate: 1/6 job/hr
Together rate: 1/x job/hr
----------------------------
Equation:
rate + rate = together rate
1/4 + 1/6 = 1/x
Multiply thru by 12x
3x + 2x = 12
5x = 12
x = 12/5 = 2.4 hrs (time to do the job if Sally and John work together.
===============
Cheers,
Stan H.